# references.bib

### v2

@@ III-C Text-guided CSI Image Generation
Our choice of latent space as that of the Stable Diffusion LDM enables text-guided manipulation of generated images without requiring an image encoding step. This capability allows us to refine or alter the visual content of CSI-generated latent representations through natural language prompts. As illustrated in Table IV, LatentCSI can generate stylistic and semantic variations of an image from the same CSI input using different text prompts. For the results shown, we employ 100 denoising steps of DDIM [song_denoising_2020] with a noise strength of 0.6 which determines the variance of noise added.
Due to the dimensionality bottleneck in LatentCSI, high-frequency details—such as facial features, fine textures, or text—are not preserved. Text-based conditioning allows us to reconstruct or augment these elements according to a specified prompt, enabling stylistic transformation and controlled editing of semantically relevant content.
@@ IV Conclusion
We presented LatentCSI, a novel approach for high-resolution image generation from CSI by leveraging the power of a pretrained latent diffusion model. Rather than generating images directly in pixel space—which is computationally expensive and often requires complex training procedures—we propose a framework that maps CSI amplitudes into the latent space of the LDM using a lightweight neural network. This design reduces computational load and simplifies training while improving perceptual quality compared to conventional end-to-end architectures, even when the end-to-end model achieves better pixel accuracy. The use of the LDM’s latent space provides a natural process to harness the power of text-guided image editing using diffusion. This allows for fine-grained control over the output, which is particularly valuable in privacy-sensitive contexts. We suggest that LatentCSI is a compelling alternative to existing CSI-to-image methods, offering superior image quality, efficiency, and controllability.
[…]

### v3

@@ III-C Text-guided CSI Image Generation
Our choice of latent space as that of the Stable Diffusion LDM enables text-guided manipulation of generated images without requiring an image encoding step. This capability allows us to refine or alter the visual content of CSI-generated latent representations through natural language prompts. As illustrated in Table IV, n generate stylistic and semantic variations of an image from the same CSI input using different text prompts. For the results shown, we employ 100 denoising steps of DDIM [13] with a noise strength of 0.6 which determines the variance of noise added.
Due to the dimensionality bottleneck in igh-frequency details—such as facial features, fine textures, or text—are not preserved. Text-based conditioning allows us to reconstruct or augment these elements according to a specified prompt, enabling stylistic transformation and controlled editing of semantically relevant content.
@@ IV Conclusion
We presented novel approach for high-resolution image generation from CSI by leveraging the power of a pretrained latent diffusion model. Rather than generating images directly in pixel space—which is computationally expensive and often requires complex training procedures—we propose a framework that maps CSI amplitudes into the latent space of the LDM using a lightweight neural network. This design reduces computational load and simplifies training while improving perceptual quality compared to conventional end-to-end architectures, even when the end-to-end model achieves better pixel accuracy. The use of the LDM’s latent space provides a natural process to harness the power of text-guided image editing using diffusion. This allows for fine-grained control over the output, which is particularly valuable in privacy-sensitive contexts. We suggest that a compelling alternative to existing CSI-to-image methods, offering superior image quality, efficiency, and controllability.